please help me solve this problem.
y=x2+4x
thats x squared.
i need it written in vertex form.
: please help me solve this problem.
y=x2+4x
thats x squared.
i need it written in vertex form.

Answer by jim_thompson5910(9869)

In order to write x^2+4x in vertex form, we need to complete the square for x^2+4x


---------------------------------------------

Completing the Square:


x^2+4x Start with the given expression.


Take half of the x coefficient 4 to get 2. In other words, (1/2)(4)=2.


Now square 2 to get 4. In other words, (2)^2=(2)(2)=4


x^2+4x+highlight(4-4) Now add and subtract 4. Notice how 4-4=0. So the expression is not changed.


(x^2+4x+4)-4 Group the first three terms.


(x+2)^2-4 Factor x^2+4x+4 to get (x+2)^2.


So after completing the square, x^2+4x transforms to (x+2)^2-4. So x^2+4x=(x+2)^2-4.



So y=x^2+4x is equivalent to y=(x+2)^2-4.


---------------------------------------------


y=(x+2)^2-4 Start with the previous equation


y=(x-(-2))^2-4 Rewrite x+2 as x-(-2)


y=1(x-(-2))^2-4 Place a "1" outside the parenthesis.


Now y=1(x-(-2))^2-4 is in vertex form y=a(x-h)^2+k where a=1, h=-2 and k=-4